# SmallPlan: Leverage Small Language Models for Sequential Path Planning with Simulation-Powered, LLM-Guided Distillation

### v3

@@ Abstract
Efficient path planning in robotics, particularly within large-scale, dynamic environments, remains a significant hurdle. While Large Language Models (LLMs) offer strong reasoning capabilities, their high computational cost and limited adaptability in dynamic scenarios hinder real-time deployment on edge devices. We present SmallPlan- a novel framework leveraging LLMs as teacher models to train lightweight Small Language Models (SLMs) for high-level path planning tasks. In SmallPlan, the SLMs provide optimal action sequences to navigate across scene graphs that compactly represent full-scaled 3D scenes. The SLMs are trained in a simulation-powered, interleaved manner with LLM-guided supervised fine-tuning (SFT) and reinforcement learning (RL). This strategy not only enables SLMs to successfully complete navigation tasks but also makes them aware of important factors like travel distance and number of trials. Through experiments, we demonstrate that the fine-tuned SLMs perform competitively with larger models like GPT-4o on sequential path planning, without suffering from hallucination and overfitting. SmallPlan is resource-efficient, making it well-suited for edge-device deployment and advancing practical autonomous robotics. Our source code is available here: https://github.com/quangpham2006/SmallPlan
Keywords: Path Planning, Distillation, Reinforcement Learning, SLMs, LLMs
Figure 1: LLMs are powerful but often costly, SLMs are efficient but lack reasoning. We bridge the gap—training SLMs for sequential path planning leveraging LLMs, simulation, SFT, and RL.
1 Introduction
@@
Recent advances in LLMs have played a key role in pushing the boundaries of robotic fields, particularly in scene understanding and path planning, due to their capability in semantic reasoning and problem-solving. [1, 2, 3] works on foundational models that utilize LLMs for common-sense reasoning and task planning. Leveraging these works, [4, 5] propose Vision-Language Models (VLMs) for long-horizon planning tasks, aiming to enable robots to perceive and traverse large-scale and complex environments. More recent studies [6, 7] explore LLMs to generate long-horizon navigation plans from topological maps [8, 9] or graph-based scene representations (scene graph) [10, 11, 12, 13]. Compared to frame-by-frame reasoning with full-scale 3D representations, these approaches offer a more lightweight and compact alternative while retaining critical global semantic and spatial context.
Nevertheless, the above-mentioned works use either heavy-weight LLMs - which are computationally intensive, or LLM-as-a-Services (GPT-4o, GPT-4) - which incur high API costs. Such setups are unsuitable for deployment on edge devices and raise concerns about increased carbon emissions [14]. Moreover, most current approaches use off-the-shelf LLMs for direct inference without retraining for fine-tuning, infeasible to adapt to changing environments or intricate task demands. Consequently, this leads to inefficient re-planning and frequently results in inexecutable action sequences [6]. Interestingly, while knowledge distillation [15] is popular in many machine learning domains including LLMs compression [16], its application and effectiveness in scene understanding and path planning remain unexplored.
@@
Therefore, we propose SmallPlan, a distillation-based framework that uses LLMs as teacher models to guide the SLMs for high-level path planning. Our framework enhances the outputs of SLMs by incorporating three key components. Firstly, the candidate SLM learns not only from the teacher LLM but also from real-time simulation outputs during fine-tuning, achieving a balance between following a stronger model and adapting to surrounding environments. Secondly, we design a custom reward policy that incentivizes the SLM to control the travel distance and number of trials, making the inference efficient and less resource-consuming using RL.
In summary, our key contributions include:
@@
  • We propose SmallPlan to enhance SLMs for sequential high-level path planning by combining LLM-guided distillation with real-time simulation-powered SFT and RL.
@@
  • We show that SLMs can achieve competitive performance with LLMs despite significantly lower resource requirements. Moreover, with customized reward in RL, SmallPlan can control side factors to reduce energy consumption, enhancing real-world deployability.
  • We experiment different training strategies and survey the contributions of SFT and RL to the fine-tuned SLMs. We analyze the critical role of SFT in stabilizing the SLMs’ outputs, and RL’s strong generalization capability in unseen environments.
2 Related Works
@@
Definition of SLMs. Categorizations of SLMs vary in their contexts [26]. For examples, [27] defines SLMs as models with fewer than 1 billion params, while [28, 29, 30] consider the term SLMs relative to the larger counterparts. For mobile devices, typically possessing around 6GB of memory, [31] suggests that SLMs contain fewer than 10 billion params. In this work, all employed SLMs are capped at 3 billion params and require no more than 4GB of memory for inference.
@@
Distillation for SLMs. Multiple studies [32, 33, 34, 35, 36] have explored techniques for transferring knowledge from large models (LLMs) to smaller ones (SLMs). One common strategy is feedback-driven fine-tuning [37, 38]. Other works [39, 40] suggest that feedback from mistakes and environment can improve fine-tuning in robotics task planning. To further enhance the fine-tuning quality, we propose a multi-objective approach to optimize path planning of SLMs. We additionally consider the SLMs’ behaviors in simulation environment to reward the model during training.
@@
SFT and RL for Fine-tuning Language Models. Numerous works have investigated how SFT and RL enhance the reasoning capability of language models. An interesting study [41] suggests that SFT tends to lead models to memorize the training data, whereas RL, especially when guided by outcome-based rewards, enables models to generalize across both rule-based textual and visual variants, thus enhancing the language model’s outcomes. Notably, the study highlights that SFT remains essential for stabilizing the model’s output format, thereby supporting reliable and effective RL training. [42] specifically surveys what works and what does not for RL for reasoning in SLMs. Inspired by these findings, our work proposes a fine-tuning strategy that utilize SFT interleaved with reward-customized RL to stabilize and enhance the outcomes of the SLMs for path planning.
3 Problem Statement and Scenario Set-up
We examine a scenario where an embodied robotic agent $\mathcal{M}$ operates within a large, unexplored environment Env. $\mathcal{M}$’s objective is to locate a specific target object $g\in\mathcal{O}$, where $\mathcal{O}$ represents the complete set of objects within Env. To accomplish this, $\mathcal{M}$ utilizes language model (SLM) for reasoning about Env and proposes sequential navigation actions toward $g$. Assumptions and constraints are defined as follows:
@@ 3 Problem Statement and Scenario Set-up
Environment Dynamics. The environment Env consists of two disjoint sets of objects: $\mathcal{O}_{\text{seen}}$, representing objects currently within the visible field of view of the robot $\mathcal{R}$, and $\mathcal{O}_{\text{unseen}}$, denoting objects outside of this view. These sets satisfy $\mathcal{O}=\mathcal{O}_{\text{seen}}\cup\mathcal{O}_{\text{unseen}}$ and $\mathcal{O}_{\text{seen}}\cap\mathcal{O}_{\text{unseen}}=\emptyset$. Transitions from $\mathcal{O}_{\text{unseen}}$ to $\mathcal{O}_{\text{seen}}$ occur through real-time perception using image or point cloud data. In the simulation environment, visibility is determined directly by the simulator engine, which updates the sets whenever an object enters the field of view of $\mathcal{M}$.
@@
4 Proposed Method
Figure 2: SmallPlan fine-tuning framework. First, an image sequence generates the hierarchical scene graph containing different rooms (Observations to text). The scene graph is fed to LLM and SLM to provide reasoning and propose next actions (High-Level Path Planning). The robot executes the action, gets the observation, and reinforces the SLM given the reward (Environment).
Fig. 2 demonstrates SmallPlan’s pipeline. At first, robot perception is fed into a model that contains a two-level hierarchical graph representing the scenes, with higher level represents a room, and lower level represents the objects located within that corresponding room. In simulation setup, information of the scenes and the outcomes of the proposed action sequences by the SLMs are extracted directly from the simulation engine. We apply chain-of-thought prompting [18] similarly to [17] to generate responses for both the LLM and the SLM. The LLM and SLM responses, together with outcomes from the simulation environment, are utilized for the fine-tuning process with SFT and RL, with details provided in Section 4.2.
4.1 From Perception to Language Model
@@
4.2 LLM-guided Distillation with Simulation Feedback
Our fine-tuning process starts by applying SFT alone to the candidate SLM for the first few epochs to stabilize its response format. Pre-trained SLMs often fail to consistently produce outputs in the required format, even when given correct prompts. This leads to problems when converting their generated responses into action commands. State-of-the-art LLMs like GPT-4o are much more reliable at following instructions and formatting responses correctly. Therefore, this initial phase helps the candidate SLM learn the correct response structure by imitating the teacher LLM.
Following this, the candidate SLM undergoes fine-tuning with the RL-interleaving-SFT approach in subsequent epochs. Each epoch completes when the task is either successfully executed or terminated early by the LLM if a deadlock situation is detected. To further accelerate the fine-tuning process, we employ Low-Rank Adaptation (LoRA), which freezes the original model weights and trains only small, additional layers instead. The fine-tuned SLM is then used for inference.
Supervised Fine-tuning.
Assuming that the teacher LLM possesses superior reasoning and planning capabilities, we treat its response as the supervision label for the candidate SLM to follow. We query both the teacher LLM and the candidate SLM for action predictions. We adopt the standard Cross-Entropy Loss as the Supervised Loss ($\mathcal{L}_{S}$) for aligning the outputs of the candidate SLM and the teacher LLM:
@@ 4.2 LLM-guided Distillation with Simulation Feedback
Reinforcement Learning with Simulation Feedback.
We reinforce the candidate SLM with Proximal Policy Optimization (PPO) algorithm with customized reward function, defined as:
@@
where $a_{t}$ denotes the action taken at time step $t$, $s_{t-1}$ is the environment state prior to the action and $s_{t}$ is the updated environment state after the action execution. A bonus reward $r_{\text{success}}$ is granted if the task is successfully completed. The task is considered successful when the action selected by the SLM is $a=\texttt{done()}$(), and the resulting graph $\mathcal{G}_{\mathcal{S}(t)}$ contains the goal object $o_{G}\in(\mathcal{O}_{\text{seen}}=\mathcal{N}_{\text{objects}})\subseteq\mathcal{G}_{\mathcal{S}(t)}$. Otherwise, rewards or penalties are assigned based on the outcomes of individual sub-actions during task execution. The action-specific reward $r_{\text{action}}$:
$r_{\text{action}}(a_{t},s_{t-1},s_{t})$ $=r_{\text{action\_success}}(a_{t},s_{t-1},s_{t})+r_{\text{explore}}(a_{t},s_{t-1},s_{t})$
$+r_{\text{efficiency}}(a_{t},s_{t-1},s_{t})+r_{\text{format}}(a_{t},s_{t-1},s_{t})$
Here, $r_{\text{action\_success}}$ provides a reward when the action is executable in the environment. $r_{\text{explore}}$ encourages the robot to explore new areas, which is essential for locating the target object. $r_{\text{efficiency}}$ penalizes excessive movement to promote efficient navigation and task execution. Finally, $r_{\text{format}}$ imposes a penalty if the response generated by the SLM does not follow the required output format. Each component of the reward function is defined as:
  • Action Success:
    $r_{\text{action\_success}}(a_{t},s_{t-1},s_{t})=\begin{cases}\lambda_{\text{success}},&\text{if the action is successfully executed,}\\ -\lambda_{\text{success}},&\text{otherwise.}\end{cases}$
@@
    $r_{\text{format}}(a_{t},s_{t-1},s_{t})=\begin{cases}0,&\text{if the action is parsed successfully,}\\ -\lambda_{\text{format}},&\text{otherwise.}\end{cases}$
    To enforce strict adherence to a predefined action format, a penalty is applied if the SLM’s output cannot be parsed into a valid executable command.
where $\lambda_{\text{action\_success}},\lambda_{\text{explore}},\lambda_{\text{efficiency}},\lambda_{\text{format}}$ are hyperparameters to balance the importance of each reward component.
5 Experiments
Our study explores the following key points: (1) The impact of SFT and RL on model performance in sequential path planning tasks; (2) trade-offs when using RL to control side factors like distance and trials; (3) the comparative effectiveness of RL versus SFT in unseen scenes; (4) the role of SFT during pre-training in enhancing subsequent RL fine-tuning; and (5) How the fine-tuned SLMs generalize to out-of-distribution tasks.
5.1 Experiment Setup
Model Selection. We investigate Qwen-2.5-3B (Qw2.5-3B)  [49] and Phi4-mini [50], each consuming approximately 4GB memory for inference with LoRA and 4-bit quantization. We also provide GPT-4o’s performance, which is used in article MoMaLLM [17], for comparison.
Simulation Setup and Dataset. We conduct experiments using the iGibson simulation platform [51, 52] with the Fetch robot, adopting the setup from [17]. The environment is configured based on the iGibson Challenge data split with 8 train scenes for development and 7 test scenes for evaluation. Each run within a scene randomizes the start poses, target objects, and object distributions, resulting in a diverse spectrum of task scenarios and action sequences. We evaluate our method over 25 run trials per test scene, yielding a total of 175 challenging and previously unseen task executions.
Out-of-domain Reasoning Analysis Setup. We conduct an out-of-domain experiment to assess the risk of overfitting in our fine-tuned SLMs. We use a text-based interactive game benchmark [53], designed to evaluate various logical reasoning skills of language models. Specifically, we compare the performance of our fine-tuned SLMs with original pre-trained LLMs and SLMs on 4 different games introduced in this benchmark.
Metrics. We provide Task Success Rate (SR %) - number of successful task executions over total trials. Success weighted by Path Length (SPL %) - fraction of distance traveled to the shortest possible path and weights it by whether the episode was successful. This metric does not consider the costs of object interactions. Distance Travel (Dist.) - length of the path traveled by the agent across all task executions. Number of Trials (Trials) - Average trial attempts by the agent per run per scene. Number of Proposed Actions (Acts.) - Average number of actions proposed by the SLM per task, including the number of high-level steps taken and the object interaction proposed.
5.2 Results

### v5

# SmallPlan: Leverage Small Language Models for Sequential Path Planning with LLM-Guided Distillation and Reinforcement Learning from Simulation Feedback

@@ Abstract
Efficient path planning in robotics, particularly within large-scale, complex environments, remains a significant hurdle. While Large Language Models (LLMs) offer strong reasoning capabilities, their high computational cost and limited adaptability hinder real-time deployment on edge devices. We present SmallPlan - a novel framework leveraging LLMs as teacher models to train lightweight Small Language Models (SLMs) for high-level path planning tasks. In SmallPlan, the SLMs provide optimal action sequences to navigate across scene graphs that compactly represent full-scaled 3D scenes. The SLMs are trained in a simulation-powered, interleaved manner with LLM-guided supervised fine-tuning (SFT) and reinforcement learning (RL). This strategy not only enables SLMs to successfully complete navigation tasks but also makes them aware of important factors like distance travel, providing more efficient path planning. Through experiments, we demonstrate that the fine-tuned SLMs perform competitively with larger models like GPT-4o on sequential path planning, without suffering from hallucination and overfitting. SmallPlan is resource-efficient, making it well-suited for edge-device deployment and advancing practical autonomous robotics. Our source code is available here: https://github.com/quangpham2006/SmallPlan
I Introduction
@@
Recent advances in LLMs have played a key role in pushing the boundaries of robotic fields, particularly in scene understanding and path planning, due to their capability in semantic reasoning and problem-solving. [1, 2, 3] works on foundational models that utilize LLMs for common-sense reasoning and task planning. Leveraging these works, [4, 5] propose Vision-Language Models for long-horizon planning tasks, aiming to enable robots to perceive and traverse large-scale and complex environments. More recent studies [6, 7] explore LLMs to generate long-horizon navigation plans from topological maps [8, 9] or graph-based scene representations (scene graph) [10, 11, 12, 13]. Compared to frame-by-frame reasoning with full-scale 3D representations, these approaches offer a lighter and compact alternative while retaining critical global semantic and spatial context.
Figure 1: LLMs are powerful but often costly, SLMs are efficient but lack reasoning. We bridge the gap—training SLMs for sequential path planning leveraging LLMs, scene graph, SFT, and RL.
However, these approaches still rely either on heavy-weight LLMs, which are computationally expensive, or on LLM-as-a-Service (GPT-4o, GPT-4) [6, 14], which incur high API costs. Such setups are impractical for edge deployment and raise concerns about carbon emissions [15]. In addition, using off-the-shelf LLMs for direct inference makes it difficult to control important factors such as the distance or cost of the planning sequences they generate, which are crucial for managing energy consumption on edge devices.
@@
Therefore, we propose SmallPlan, a distillation framework that uses LLMs as teacher models to train the SLMs for high-level path planning. Our framework enhances the outputs of SLMs by incorporating two key components. Firstly, the candidate SLM learns not only from the teacher LLM but also from real-time simulation feedback during training, achieving a balance between following a stronger model and adapting to surrounding environments. Secondly, we leverage RL and design a custom reward policy that incentivizes the SLM to control the travel distance and number of trials, making the inference less resource-consuming.
Our key contributions, demonstrated in Fig. 1, include:
@@
  • We enhance SLMs for sequential high-level path planning by combining LLM-guided distillation fine-tuning and RL from simulation feedback.
@@
  • We show that SLMs can achieve competitive performance with LLMs despite significantly lower resource requirements. SmallPlan enhances the SLMs’ capability by learning from superior LLMs. Moreover, with RL and our reward policy, SmallPlan controls important factors to make the planning more efficient, enhancing real-world deployability.
  • We study different training strategies that contribute to the outcome SLMs. We analyze the critical role of SFT distillation in stabilizing the SLMs’ outputs, and how the SLMs adapt to different reward penalties in RL.
II Related Works
@@
Definition of SLMs. Categorizations of SLMs vary in their contexts [24]. For examples, [25] defines SLMs as models with fewer than 1 billion params, while [26, 27, 28] consider the term SLMs relative to the larger counterparts. For mobile devices, typically possessing around 6GB of memory, [29] suggests that SLMs contain fewer than 10 billion params. In this work, all employed SLMs are capped at 3 billion params and require less than 4GB memory for inference.
@@
Distillation for SLMs. Multiple studies [30, 31, 32, 33, 34] have explored techniques for transferring knowledge from large models (LLMs) to smaller ones (SLMs). One common strategy is feedback-driven fine-tuning [35]. Other works [36, 37] suggest that feedback from mistakes and environment can improve fine-tuning in robotics task planning. [38] proposes a SFT distillation framework for turning point robot navigation. To further explore the capability of SLMs, we propose a multi-objective approach by additionally consider the SLMs’ behaviors in simulation environment to reward the model with RL.
@@
SFT and RL for Fine-tuning [^1] In our paper, ”fine-tuning” and ”training” are using interchangeably Language Models. Numerous works have investigated how SFT and RL enhance the reasoning capability of language models. An interesting study [39] suggests that SFT tends to lead models to memorize the training data, whereas RL, especially when guided by outcome-based rewards, enables models to generalize across both rule-based textual and visual variants, thus enhancing the language model’s outcomes. Notably, the study highlights that SFT remains essential for stabilizing the model’s output format, thereby supporting reliable and effective RL training. [40] specifically surveys what works and what does not for RL for reasoning in SLMs. Inspired by these findings, our work proposes a fine-tuning strategy that utilize SFT interleaved with reward-customized RL to stabilize and enhance the outcomes of the SLMs for path planning.
Figure 2: SmallPlan training framework. (Perception) A hierarchical scene graph constructed from robot observations, representing rooms and their associated objects. (Observations to Text) The scene graph then converted into text and feed to language models for path planning. (High-Level Path Planning) The LLM and SLM use the graph as context to propose the next actions. (Environment) The robot executes actions, receives new observations, and updates the SLM using rewards.
III Problem Statement and Scenario Set-up
We consider a scenario in which an embodied robotic agent $\mathcal{M}$ operates in a large, previously unexplored environment Env. The agent’s objective is to locate a target object $o_{g}\in\mathcal{O}$, where $\mathcal{O}$ denotes the complete set of objects within Env. To achieve this, $\mathcal{M}$ leverages a language model (SLM) to reason about Env and generate sequential navigation actions that guide it toward $g$. Here, sequential navigation refers to a sequence of correlated actions drawn from an available action pool, where each chosen action builds upon the previous ones. This enables $\mathcal{M}$ to interact with both objects and the environment in a coherent manner to successfully complete the task. The following assumptions and constraints are imposed:
@@ III Problem Statement and Scenario Set-up
Environment Scenario. The environment Env consists of two disjoint sets of objects: $\mathcal{O}_{\text{seen}}$, representing objects currently within the visible field of view of the robot $\mathcal{R}$, and $\mathcal{O}_{\text{unseen}}$, denoting objects outside of this view. These sets satisfy $\mathcal{O}=\mathcal{O}_{\text{seen}}\cup\mathcal{O}_{\text{unseen}}$ and $\mathcal{O}_{\text{seen}}\cap\mathcal{O}_{\text{unseen}}=\emptyset$. Transitions from $\mathcal{O}_{\text{unseen}}$ to $\mathcal{O}_{\text{seen}}$ occur through real-time perception using image or point cloud data. In the simulation environment, visibility is determined directly by the simulator engine, which updates the sets whenever an object enters the field of view of $\mathcal{M}$.
@@
IV Proposed Method
Fig. 2 demonstrates SmallPlan’s pipeline. At first, robot perception is fed into a model that contains a two-level hierarchical graph representing the scenes, with higher level represents a room, and lower level represents the objects located within that corresponding room. In simulation setup, information of the scenes and the outcomes of the proposed action sequences by the SLMs are extracted directly from the simulation engine. We apply chain-of-thought prompting [16] similarly to [14] to generate responses for both the LLM and the SLM. The LLM and SLM responses, together with outcomes from the simulation environment, are utilized for the fine-tuning process with SFT and RL, with details provided in Section IV-B.
IV-A From Perception to Language Model
@@
IV-B LLM-guided Distillation and RL with Simulation Feedback
IV-B1 Supervised Fine-tuning with teacher LLM
Assuming that the teacher LLM possesses superior reasoning and planning capabilities, we treat its response as the supervision label for the candidate SLM to follow. This is the distillation. We query both the teacher LLM and the candidate SLM for action predictions. We adopt the standard Cross-Entropy Loss as the Supervised Loss ($\mathcal{L}_{S}$) for aligning the outputs of the candidate SLM and the teacher LLM:
@@ IV-B1 Supervised Fine-tuning with teacher LLM
IV-B2 Reinforcement Learning with Simulation Feedback
We reinforce the candidate SLM with the Proximal Policy Optimization (PPO) algorithm with a customized reward function, defined as:
@@ IV-B2 Reinforcement Learning with Simulation Feedback
where $a_{t}$ denotes the action taken at time step $t$, $s_{t-1}$ is the environment state prior to the action and $s_{t}$ is the updated environment state after the action execution. A bonus reward $r_{\text{success}}$ is granted if the task is successfully completed. The task is considered successful when the action selected by the SLM is $a=\texttt{done()}$, and the resulting graph $\mathcal{G}_{\mathcal{S}(t)}$ contains the goal object $o_{g}\in(\mathcal{O}_{\text{seen}}=\mathcal{N}_{\text{objects}})\subseteq\mathcal{G}_{\mathcal{S}(t)}$. Otherwise, rewards or penalties are assigned based on the outcomes of individual sub-actions during task execution. The action-specific reward $r_{\text{action}}$:
$r_{\text{action}}(a_{t},s_{t-1},s_{t})$ $=r_{\text{action\_executable}}(a_{t},s_{t-1},s_{t})$
$+r_{\text{explore}}(a_{t},s_{t-1},s_{t})$
$+r_{\text{efficiency}}(a_{t},s_{t-1},s_{t})$
$+r_{\text{format}}(a_{t},s_{t-1},s_{t})$
Here, $r_{\text{action\_executable}}$ provides a reward when the action is executable in the environment. $r_{\text{explore}}$ encourages the robot to explore new areas, which is essential for locating the target object. $r_{\text{efficiency}}$ penalizes excessive movement to promote efficient navigation and task execution. Finally, $r_{\text{format}}$ imposes a penalty if the response generated by the SLM does not follow the required output format. Each component of the reward function is defined as:
  • Action Executable:
    $r_{\text{action\_executable}}(a_{t},s_{t-1},s_{t})=\begin{cases}\lambda_{\text{executable}},\ \text{if the action is}\\ \ \ \text{successfully executed,}\\ -\lambda_{\text{executable}},\ \text{otherwise.}\end{cases}$
@@
    $r_{\text{format}}(a_{t},s_{t-1},s_{t})=\begin{cases}0,\text{if the action is parsed}\\ \text{ successfully,}\\ -\lambda_{\text{format}},\ \text{otherwise.}\end{cases}$
    A penalty is applied if the SLM response cannot be parsed into a valid executable command.
where $\lambda_{\text{action\_executable}},\lambda_{\text{explore}},\lambda_{\text{efficiency}},\lambda_{\text{format}}$ are hyper-params that define the importance of each reward component.
TABLE I: SmallPlan’s SLMs, trained with RL-SFT strategy, across 7 test scenes with 25 runs per scene, randomized start poses, targets, and object distributions.
Baseline Model SR % $\uparrow$ SPL % $\uparrow$ Dist. $\downarrow$ Retrials $\downarrow$
MoMaLLM [14] GPT-4o 97.14 65.92 18.01 0.39
Pretrained Qwen-2.5 61.33 46.59 12.00 3.05
Pretrained Llama-3 53.68 38.02 9.84 4.16
Pretrained Phi4-mini 70.86 41.57 17.84 3.39
Ours Qwen-2.5 83.43 48.07 19.87 1.04
Ours Llama-3 93.14 66.10 17.65 1.22
Ours Phi4-mini 96.57 70.41 16.13 0.49
IV-B3 SmallPlan Training Strategy
Algorithm 1 provides the pseudo-code for SmallPlan’s training procedure, consisting of 2 stages. We employ Low-Rank Adaptation (LoRA) to accelerate the training process, which freezes the original model weights and focuses on modifying a smaller subset of weights to reduce computational and memory overhead.
Algorithm 1 SmallPlan Training Strategy
1:prompt_templateprompt\_template, LLMLLM, SLMSLM, num_epochsnum\_epochs, num_epochs_fewshotnum\_epochs\_fewshot, taskstasks
2:Stage 1: Few-shot SFT for Format Refinement
3:for $epoch\leftarrow 1$ to num_epochs_fewshotnum\_epochs\_fewshot do